OpenStudy (anonymous):

Really need help solving this problem. A thousand points are graphed in a coordinate plane. Explain why it is possible to draw a straight line in the plane so that half of the points are on one side of the line and half are on the other. (Consider the slopes of the lines determined by each pair of points.)

OpenStudy (anonymous):

Really need help solving this step by step. I know the answer is 2 but i dont know how to get the answer by doing it on paper\[\frac{\sqrt{2}+\sqrt{6}}{\sqrt{2+\sqrt{3}}}\]

OpenStudy (anonymous):

you can easily draw a straight line through the pts by finding semi averages......the standard deviation of both sides of the line must be the same if it the line of best fit.......remember if it is a liner relationship it is possible to draw a straight line the fluctuations in pts might be due to experimental errors by drawing the line of best fit u are actually trying to rule out these errors

OpenStudy (anonymous):

so would you only be able to do this with linear equations?

OpenStudy (anonymous):

Let \[y = \frac{ \sqrt{2} + \sqrt{6}}{\sqrt{2 + \sqrt{3}\]

OpenStudy (anonymous):

oops !

OpenStudy (anonymous):

\[y = \frac{\sqrt2 + \sqrt3}{\sqrt{2 +\sqrt3}}\]

OpenStudy (anonymous):

then square them

OpenStudy (anonymous):

oh no typo again its 6 not 3 in numerator

OpenStudy (anonymous):

\[y^2 = \frac{ 2 + 6 - 2 \times 2 \times \sqrt3}{2 + \sqrt3}\]

OpenStudy (anonymous):

it gives \[y^2 = 4 \] \[y = \pm 2\] ...But you take y = 2

OpenStudy (anonymous):

To answer the first question ie "A thousand points are graphed in a coordinate plane. Explain why it is possible to draw a straight line in the plane so that half of the points are on one side of the line and half are on the other. (Consider the slopes of the lines determined by each pair of points.)" If u imagine an arbitrary line in the plane and then for each point make a point on the line at the perpendicular distance, then mark a point on the line such that half of the points are on one side of it, then the perpendicular line at that point will divide all the points into 2 equal halves (see pic) Might be some odd possibilities, lines in arow, that sort of thing, u can just select a different arbitrary line to get round it, the principle involved is still valid.

OpenStudy (anonymous):

ishaan94- howd did you get that the y^2=4? i typed it into the calculator and i didnt get four

OpenStudy (anonymous):

if not linear then you can make them linear by linear law

OpenStudy (anonymous):

Courtney16 top is 8 + 4 root3 = 4(2 + root 3) so 4. (ishaan94 made a typo with the sign)

OpenStudy (anonymous):

estudia-does this answer have anything to do with the slopes though?

OpenStudy (anonymous):

and thank you!

OpenStudy (anonymous):

oh yeah typo sorry thanks estudier

OpenStudy (anonymous):

Julia asked me the same thing:-) U can see the answer I gave to her (basically the "helpful hint" is not helpful(to me) since I can't thin of any rule to associate pairs of points ie if I choose one pair, I then have 999 candidates for the other). Doesn't mean there is not some method, I am thinking about it...

OpenStudy (anonymous):

can i see the link to see what you told her?

OpenStudy (anonymous):

Besides, I like my answer....:-)

OpenStudy (anonymous):

Not Julia (smacks head) tulip http://openstudy.com/users/tulip235/updates/4e32e0770b8ba7b2da418624

OpenStudy (anonymous):

ha its ok.. for some reason its not clicking with me!so you basically draw a line and then every point you have...i dont understand what you do to have half the points on each side

OpenStudy (anonymous):

Did u look at the pic? U have projected all the points perpendicularly (a projection) onto an arbitrary line. This gives them a linear order and u can count them. So u can select (on the arbitrary line) a point between the 2 middle points and draw a perpendicular to divide the points in two halves. This demonstrates that it can be done although I haven't considered some edge cases (u can always shift the arbitrary line)

OpenStudy (anonymous):

Try putting all of the points on one side of the arbitrary line and follow the procedure...

OpenStudy (anonymous):

ohh okay..i think!! so if you graphed a thousand points, you would just draw an arbitrary line and then draw a perpendicular line inbetween 500 and 501?

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

thank you soo much haha!! i dont understand why it took me that long to figure that out! and for that graphing question you answered for tulip,is it really just y=5 y=-x and y=x?

OpenStudy (anonymous):

Looks that way...

OpenStudy (anonymous):

Cubic in y so 3 roots.

OpenStudy (anonymous):

ok! and i just saw another person ask the exact same question as what i did haha!

OpenStudy (anonymous):

\[\frac{ \sqrt{2}+\sqrt{6} }{ \sqrt{2+\sqrt{3}} }=\frac{ \sqrt{2}+\sqrt{6} }{ \sqrt{\frac{ 1 }{ 4 }\left( 8+4\sqrt{3} \right)} }=\frac{ 2\left( \sqrt{2}+\sqrt{6} \right) }{ \sqrt{2+2\sqrt{12}+6} }=\frac{ 2\left( \sqrt{2}+\sqrt{6} \right) }{ \sqrt{\left( \sqrt{2}+\sqrt{6} \right)^{2}} }=\frac{ 2\left( \sqrt{2}+\sqrt{6} \right) }{ \sqrt{2}+\sqrt{6} }=2\] This is the correct approach to the problem that was given, using the Perfect Square Special Factoring Formula.